OpenStudy (anonymous):

What is the first five terms in the geometric sequence An=(2)n-1

OpenStudy (amoodarya):

geometric sequence ? do you mean a(n)=2^(n-1)

OpenStudy (anonymous):

Yes

OpenStudy (amoodarya):

put n=1 ,2, 3,4,5 in it like this a(2)=2^(2-1)=2^1=2 a(1)=2^(1-1)=2^0=1